Hello People,
Trying to figure out and understand graphs, here is my question.
Please Help! I working on a two-step process. Thanking you in advance.
SEE BELOW::::
Write the equation for this situation: the value, y, in hundreds of dollars, of Allen’s stamp collection, is 2 dollars more than 2/3 of the age, x (in years).
Then graph the equation and use the graph to determine the value of Allen’s coin collection if it is 3 years old.
Betty

Answer by Theo(13342)

y = 2/3 * x + 2


if x = 0, then y = 2/3 * 0 + 2 = 2
if x = 1, then y = 2/3 * 1 + 2 = 2/3 + 2 = 2/3 + 6/3 = 8/3 = 2 and 2/3
if x = 2, then y = 2/3 * 2 + 2 = 4/3 + 2 = 4/3 + 6/3 = 10/3 = 3 and 1/3
if x = 3, then y = 2/3 * 3 + 2 = 6/3 + 2 = 6/3 + 6/3 = 12/3 = 4

the graph of the equation looks like t his:

$$$

when allen's collection is 3 years old, it's value of 400 dollars.

that's because the value of y is in hundreds of dollars.

you could also have graphed this as:

y = 200/3 * x + 200, in which case the value of y would have been in dollars.

the graph of that equation is shown below:

$$$

in order to see that, you needed to change the scale of your y-axis from each unit representing 1 to each unit representing 100.